# Joint2Human: High-quality 3D Human Generation via Compact Spherical Embedding of 3D Joints

### 14 Dec 2023 v1

@@ Abstract
3D human generation is increasingly significant in various applications. However, the direct use of 2D generative methods in 3D generation often results in significant loss of local details, while methods that reconstruct geometry from generated images struggle with global view consistency. In this work, we introduce Joint2Human, a novel method that leverages 2D diffusion models to generate detailed 3D human geometry directly, ensuring both global structure and local details. To achieve this, we employ the Fourier occupancy field (FOF) representation, enabling the direct production of 3D shapes as preliminary results using 2D generative models. With the proposed high-frequency enhancer and the multi-view recarving strategy, our method can seamlessly integrate the details from different views into a uniform global shape. To better utilize the 3D human prior and enhance control over the generated geometry, we introduce a compact spherical embedding of 3D joints. This allows for effective application of pose guidance during the generation process. Additionally, our method is capable of generating 3D humans guided by textual inputs. Our experimental results demonstrate the capability of our method to ensure global structure, local details, high resolution, and low computational cost, simultaneously. More results and code can be found on our project page at http://cic.tju.edu.cn/faculty/likun/projects/Joint2Human.
Figure 1: Our method can generate high-quality and high-resolution 3D humans with reasonable global structure and fine-grained geometry details in an efficient way, based on our proposed compact spherical embedding of 3D joints and a direct conditional 3D generative network with 2D diffusion models.
@@ 1 Introduction
The field of 3D human generation holds considerable significance across diverse applications, including virtual/augmented reality, gaming, and the film industry. Despite advancements in existing methods [21, 9, 44], achieving a simultaneous guarantee of both global structural accuracy and preservation of local details remains a challenge, often compounded by high computational costs. In this paper, we aim to address these limitations and propose a methodology for generating high-quality 3D human models that exhibit fidelity to both global structures and local details, while ensuring computational efficiency, as illustrated in Figure 1.
Existing human generation methods can be divided into two categories: two-step generation methods and direct 3D generation methods. Two-step generation methods [18, 19, 44, 21, 32] generate 3D humans from 2D images via Nerf [28] or differentiable rendering [29]. These methods are trained only on 2D datasets and hence lack the perception of 3D global structures, which results in view inconsistency. Chupa [21] aims to produce high-quality results through dual normal map-based optimization but suffers from depth ambiguity issues that harm view consistency. Direct 3D generation methods [1, 31, 9, 42] generate 3D humans directly with Tri-planes [4] or Signed Distance Fields [32]. However, these approaches encounter challenges in achieving high-fidelity global structure with fine-grained local geometry details. Additionally, their generalization capabilities are limited, and computational efficiency is compromised. In summation, the existing methodologies collectively fall short in simultaneously ensuring global structural fidelity, local detail preservation, high resolution, and computational efficiency.
To achieve efficiently generating high-quality 3D humans with reasonable global structure and fine-grained geometry details, in this paper, we propose Joint2Human, a conditional generative network with 2D diffusion models derived from 3D datasets. To achieve high-fidelity 3D human generation with reasonable global structure, we employ an image-aligned 3D representation FOF to directly produce 3D shapes as preliminary results with 2D generative models. Additionally, we propose a compact spherical embedding of 3D joints for effective pose guidance, enhancing the overall fidelity of generated poses. We also design a high-frequency enhancer and a multi-view recarving strategy for fine-grained local detail generation. Experimental results demonstrate that our method outperforms the state-of-the-art methods in terms of global structure, local detail, and efficiency. Furthermore, our method also exhibits versatility by enabling the generation of 3D human representations guided by text.
@@
  • We propose Joint2Human, a direct conditional 3D generative method with 2D diffusion model for high-quality, high-resolution 3D human generation. To our best knowledge, our work is the first work that can ensure global structure, local details, high resolution, and low computational cost, simultaneously.
  • We propose a new 3D pose representation, a compact spherical embedding of 3D human joints, for efficient perception of global structure. This mechanism also facilitates a more straightforward and effective implementation of pose-guided generation.
  • We design a high-frequency enhancer by integrating a subsidiary decoder into the pre-trained VAE, and a multi-view recarving strategy, for fine-grained local detail generation.
@@ 2 Related Work
Many approaches [18, 19, 44, 21, 32, 11] try to learn the 3D shape from 2D images via various Nerf representations [28, 30, 25, 35, 36, 23] and differentiable volume rendering [29, 48, 43]. It is always computationally expensive and limited in resolution. EVA3d [18] is carefully designed in terms of training strategy although it achieves high-resolution generation. What’s worse, the 2D dataset is always imbalanced in viewing angles and human poses, so this strategy is relatively unstable and hard to generate a realistic full-body human geometry. Other methods  [19, 44, 21] also seek help with priors from 2D human generation and 3D human reconstruction models, which is cumbersome. In Chupa [21], many optimizations are needed to maintain viewpoint consistency. It first generates two normal maps for the front and back sides of a clothed human, then optimizes the 3D generated mesh by the dual normal maps. There are also some human template-based works  [47, 17, 45, 3, 21] that are highly dependent on human template mesh, such as SMPL [26] and SMPL-X [34], which adversely affect the diversity of generated models especially in local details. For example, it is flawed in modeling loose clothes like dresses.
@@
Based on the development of 3D human representation, numerous methods have emerged for direct 3D generation. Some of these approaches [1, 31, 9, 42] utilize tri-planes [4] which is a Nerf-based representation. It is hard to generate the precise geometry limited by the computational complexity. Apart from that, the implicit functions like SDFs [32, 33] are highly favored in several methods [44, 3, 18, 17, 22]. Alternatively, some methods use 3D data for training, gDNA [6] use implicit multi-subject forward skinning which enables learning from 3D scans of human bodies. Among all the works mentioned above, the GAN [12, 47, 18, 1, 31, 9, 19, 4, 32, 44] and Diffusion models [16, 42, 3, 21, 22] are the most popular. There are also some human template-based works  [47, 17, 45, 3, 21] that are highly dependent on human template mesh, such as SMPL [26] and SMPL-X [34], which adversely affect the diversity of generated models especially in local details. For example, it is flawed in modeling loose clothes like dresses.
Using the diffusion model for small object generation[7, 24, 8, 5] is easier to combine with text control. However, the text-guided 3D human generation is more difficult due to the high-dimensional cube and lack of text-3D data pairs. In such a situation, some attempts were made. AvatarCLIP [17] initialize a bare mesh shape with SMPL, then it continuously optimizes the mesh to match the input text by calculating the CLIP score between the text token and the rendering image of mesh, CLIP-actor [45] does the same thing except it designs a recommendation module to get the initial shape/motion. In this way, they reduced the need for text-3D data pairs with the help of the large text-image pretrained model [37]. Other methods [21, 3, 42] impose textual controls on the diffusion model for generated shapes that match the input description.
@@
3 Preliminaries
Figure 2: Overview of our framework. Joint2Human takes gaussian noise and some conditional information (pose,text) as input. To enable high-resolution human generation, we first encode FOF into a latent space, where a diffusion model is trained. Furthermore, to enable flexible conditional generation, we adopt condition encoders along with classifier-free guidance to enable conditional generation. Our conditional control generation strategy can support switching between different modalities. Then, we pass the results generated by the first denoising process through the high-frequency enhancer $\mathcal{D}_{h}$ and multi-view recarving strategy for fine-grained local detail generation. Additionaly, the recaiving strategy is shown in the bottom subfigure.
3.1 Fourier Occupancy Fields
Fourier Occupancy Field (FOF) [10] is an efficient 3D representation based on the Fourier series expansion of a square-wavelike function. It spans the 3D occupancy field along the view direction and takes the low frequency terms to represent the human geometry with a 2D map. This can be written as:
$F(x,y,z)=b^{T}(z)C(x,y),$ (1)
where $b^{T}(\cdot)$ is the basic functions, and $C(\cdot,\cdot)$ is the FOF. Such a representation enables direct inference of 3D geometry using 2D frameworks. Thanks to the convenient conversion between the FOF and 3D occupancy field, we can integrate the details from different views into a uniform 3D field with our multi-view recarving strategy, which is described in Sec. 4.3.
3.2 Diffusion Probabilistic Model
Diffusion Probabilistic Models [16] is a generative model that could realize high-quality image generation. It contains two processes, the fixed forward process $q$ and backward processes $p$ which can be viewed as a Markov chain indexed by timestep $t$. In the forward process, we gradually add Gaussian noise $\epsilon$ to raw datapoint $z_{0}$ in each timestep t, we have:
$z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,$ (2)
where we use a sequence of increasing noise schedules $\beta_{t}\in\left\{\beta_{1,}...\beta_{T}\right\}$ with the last timestep $T$, $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod\nolimits_{s=1}^{t}{\alpha_{s}}$. After sufficient noising steps, the raw data $z_{0}$ is transformed to pure Gaussian noise. Then it’s time for the backward process $p$, which adopts a U-Net [39] structure as the denoising model $\epsilon_{\theta}$, trained to predict the noise added at every step for removing noise. The loss functions can be expressed simply in the form of:
$\mathcal{L}_{DDPM}(\theta):=\mathbb{E}_{\mathbf{z},\epsilon\sim N(0,1),t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\mathbf{z}_{t},t\right)\right\|^{2}\right].$ (3)
For generation, we first sample a noise $z_{t}\sim N\left(0,1\right)$ from the Gaussian distribution during inference time. Then we gradually denoise $z_{t}$ step by step by estimating $\epsilon$. The reverse process can be formulated as:
$\mu_{\theta}\left(z_{t-1},t\right)=\frac{1}{\sqrt{\alpha_{t}}}\left(z_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}\left(z_{t},t\right)\right).$ (4)
We can obtain the final results by iteratively computing $z_{t-1}=\mu_{\theta}\left(z_{t-1},t\right)+\sigma_{t}\epsilon$ along with the denoising preocess.
@@ 4 Method
Our work aims at generating 3D full-body human models with global structure and local details. To achieve this, we design an effective pipeline as illustrated in Fig. 2. The highlight of our work is that we propose a direct conditional 3D generative method with 2D diffusion models. Specifically, we use an image-aligned 3D representation FOF [10] to directly produce 3D shapes. Then, we model the distribution of FOF features. However, the raw data space is high-dimensional and high-resolution, so modeling these feature maps is too costly to compute. So, we first map Fourier Occupancy Fields to a low dimensional space and then operate our condition and generation process in the latent space (Sec. 4.1). For reasonable global structure and effective pose guidance, we propose a compact spherical embedding of the 3D human joints (Sec. 4.2). Furthermore, We also design a high-frequency enhancer (Sec. 4.3) and a multi-view recarving strategy (Sec. 4.4) for fine-grained local detail generation. We will detail all the stages of our method in the following subsections.
@@ 4.1 Latent Diffusion for Fourier Occupancy Fields
To model the distribution of Fourier Occupancy Fields, following the latent diffusion model [38], we utilize an auto-encoder to compress the high-dimensional raw data space into a lower-dimensional latent space, which encodes each shape into a normal distribution. In detail, we adopted a VAE-like auto-encoder which contains the encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Given a FOF feature $x\in\,\,\mathbb{R}^{512\times 512\times 32}$, the encoder $\mathcal{E}$ encodes FOF into latent vectors $z=\mathcal{E}\left(x\right)$, where $z\in\mathbb{R}^{128\times 128\times 8}$. The decoder $\mathcal{D}$ decodes the latent vectors back to the original FOF space $\tilde{x}=\mathcal{D}\left(\mathcal{E}\left(x\right)\right)$, where $\tilde{x}\in\mathbb{R}^{512\times 512\times 32}$. We pretrained the auto-encoder with the reconstruction loss and KL-regularization loss. Before DDPM model training, we freeze the weights of the VAE, and the training data is passed through the encoder $\mathcal{E}$ in the vae module to obtain its feature in the latent space. We adopt a U-Net-like structure as our denoising model $\epsilon_{\theta}$. We follow the classical and efficient loss function proposed by Ho [16] and train the diffusion model based on the feature from noise with the diffusion model in the latent space. Thus we can obtain a vector $\tilde{z}_{0}$ in the latent space, which will be sent to the decoder $\mathcal{D}$ to get Fourier occupancy fields. Then, we can get the 3D occupancy field from FOF. We can get the 3D human mesh from the 3D occupancy field with the Marching Cubes [27] algorithm. In this way, we have further mitigated the computational cost of 3D representations.
Figure 3: The generation process for compact spherical embedding of 3D joints. Given a human mesh, we first compute its 3D joint points. Then we generate a solid ball around each joint and convert them to FOF. For the compact spherical embedding of 3D joints, we only retained the low-frequency information, so the visualization results are not strictly spherical.
@@ 4.2 Condition-guided Generation Mechanisms
Compact Spherical Embedding of 3D Joints. Instead of adopting the pose-guidance from the human parametric models [26, 34]. Many previous works [21, 9, 47, 45, 17, 3] rely on SMPL or SMPL-X as pose guidances, which adversely affects the diversity of generated models especially in local details. Regardless of whether SMPL mesh or a normal map from SMPL is used as the reference to generate a clothed human body, this information from the strong geometric prior is redundant. Thus, our model may be misleading to learn the extra knowledge. As a result, it may memorize the training set during training. Even during inference, the generated result closely resembles the highly desired results from the training set. To avoid denoising model $\epsilon_{\theta}$ from being in an over-fitting regime, we propose compact spherical embedding for 3D human joints, encompassing both the 2D position and depth in 3D space as shown in Fig. 3 we only save the $K$ joints location $J=\left\{p_{i}\in\mathbb{R}^{3}\right\}_{i=1}^{K}$ from the whole human body as our guidance. For each joint position $p_{i}$, We extend a solid sphere with $p_{i}$ as the center and $r$ as the radius in 3D space. After that, according to the Sec. 3.1, we compute the FOF $C_{i}\in\mathbb{R}^{512\times 512\times 8}$ of each solid sphere. Then we concat $\left\{C_{i}\right\}_{i=1}^{K}$ channel-wise as our joint condition $con_{joint}\in\mathbb{R}^{512\times 512\times 8K}$. During training, we apply the joint condition $con_{joint}$ on the latent code $z$ after passing through the conditional encoder.
@@
Pose-guided Geometry Generation. Given a 3D human pose, our goal is to generate a human body shape that fits the given pose. With the compact spherical embedding of the 3D joints, we can efficiently perform pose control of human generation. After observing the generated results, we found that we occasionally encountered the issue of inconsistent orientation between the upper and lower halves of the generated human. To address it, we incorporate 2D prior information by utilizing the IUV map as defined in DensePose [13]. In the training stage, we learn a pose encoder to get the human pose embedding $\tau_{pose}$, then apply it to the training of the diffusion model. In the inference stage, we first get the human pose embedding $\tau_{pose}$ from the input pose priori, then we perform the denoising process of the $\epsilon_{\theta}$ to get a 3D human shape.
Text-guided Geometry Generation. Given a textual description, our goal is to generate a 3D human that fits the text input. Previous methods often rely on the CLIP model to extract the embeddings from text and images, these features play an important role in the process of training and testing. However, this strategy depends on the textual annotation of the training data that is missing for us.
@@
To generate the 3D human with the guidance of text description, we proposed the image prompt strategy for our model based on a text-guided 2D human image generation model Text2Human [20]. In the training stage, We learn an additional image encoder to get the human image embedding $\tau_{img}$, then apply it to the training of the diffusion model. In the inference stage, with the input text, we first infer the text-to-image model $\mathcal{T}$ to get the human image $I_{h}$, then we sample the 3D human shape from $\epsilon_{\theta}$. We provide additional text-guided generation results in the supplementary materials.

### 06 Apr 2024 v2

@@ Abstract
3D human generation is increasingly significant in various applications. However, the direct use of 2D generative methods in 3D generation often results in losing local details, while methods that reconstruct geometry from generated images struggle with global view consistency. In this work, we introduce Joint2Human, a novel method that leverages 2D diffusion models to generate detailed 3D human geometry directly, ensuring both global structure and local details. To achieve this, we employ the Fourier occupancy field (FOF) representation, enabling the direct generation of 3D shapes as preliminary results with 2D generative models. With the proposed high-frequency enhancer and the multi-view recarving strategy, our method can seamlessly integrate the details from different views into a uniform global shape. To better utilize the 3D human prior and enhance control over the generated geometry, we introduce a compact spherical embedding of 3D joints. This allows for an effective guidance of pose during the generation process. Additionally, our method can generate 3D humans guided by textual inputs. Our experimental results demonstrate the capability of our method to ensure global structure, local details, high resolution, and low computational cost simultaneously. More results and the code can be found on our project page at http://cic.tju.edu.cn/faculty/likun/projects/Joint2Human.
Figure 1: With the proposed compact spherical embedding of 3D joints, our method can generate high-quality and high-resolution 3D humans with reasonable global structure and fine-grained geometry details in an efficient way, based on a native conditional 3D generative network with 2D diffusion model.
@@ 1 Introduction
The field of 3D human generation holds considerable significance across diverse applications, including virtual/augmented reality, gaming, and the film industry. Despite advancements in existing methods [22, 9, 46], it remains a challenge to achieve a simultaneous guarantee of both global structural accuracy and local geometry details, which is often compounded by high computational costs. In this paper, we aim to address these limitations and propose a method for generating high-quality 3D human models that exhibit fidelity to both global structures and local details while ensuring computational efficiency, as illustrated in Figure 1.
Existing human generation methods can be divided into two categories: two-stage methods and native 3D generation methods. Two-stage generation methods [19, 20, 46, 22, 35] fit 3D humans from 2D images via NeRF [31] or differentiable rendering [32]. These methods are trained only on 2D datasets and hence lack a 3D backbone for perceiving 3D structures and ensuring view consistency. Chupa [22] aims to produce high-quality results through dual normal map-based optimization but suffers from depth ambiguity issues that harm view consistency. Most native 3D generation methods [1, 34, 9, 44] generate 3D humans directly with Tri-planes [4] or Signed Distance Fields [35]. However, these approaches encounter challenges in achieving high-fidelity global structures with fine-grained local geometry details. Additionally, their generalization capabilities are limited, and computational efficiency is compromised. In summary, the existing methodologies collectively fail to ensure global structural fidelity, local detail preservation, high resolution, and computational efficiency simultaneously.
To efficiently generate high-quality 3D humans with reasonable global structure and fine-grained geometry details, in this paper, we propose Joint2Human, a conditional generative network with 2D diffusion models derived from 3D datasets. To achieve high-fidelity 3D human generation with reasonable global structure, we employ the 2D generative models to produce 3D shapes as preliminary results directly. Subsequently, we carefully design the compact spherical embedding of 3D joints based on an image-aligned 3D representation FOF [11, 10]. With it, we implement effective pose guidance and diverse generation. We also design a high-frequency enhancer and a multi-view recarving strategy for fine-grained local detail generation. Experimental results demonstrate that our method outperforms the state-of-the-art methods regarding global structure, local detail, and computational efficiency. Furthermore, our method also exhibits versatility by enabling the generation of 3D human representations guided by text.
@@
  • We propose Joint2Human, a native conditional 3D generative method with a 2D diffusion model for high-quality, high-resolution 3D human generation. To our knowledge, it is the first work based on the FOF that can simultaneously ensure global structure, local details, high resolution, and low computational cost.
  • We propose a new pose guidance embedding, a compact spherical embedding of 3D human joints, for efficient perception of global structure. This mechanism also facilitates a more straightforward and effective implementation of pose-guided generation in 2D generation framework.
  • We design a high-frequency enhancer by integrating a subsidiary decoder into the pre-trained VAE and a multi-view recarving strategy for fine-grained local detail generation. Both of them improve the geometry quality of the final results.
@@ 2 Related Work
Many approaches [19, 20, 46, 22, 35, 12] try to learn the 3D shape from 2D images via various NeRF representations [31, 33, 28, 38, 39, 25] and differentiable volume rendering [32, 51, 45]. However, it is always computationally expensive and limited in resolution. EVA3d [19] designs complex training strategies to achieve high-resolution generation, but its geometric quality is mediocre. What’s worse, the 2D dataset is always imbalanced in viewing angles and human poses. Hence, these methods are relatively unstable, and it is hard to generate a realistic full-body human geometry. Other methods  [20, 46, 22] also leverage the priors from 2D human generation and 3D human reconstruction models, which is cumbersome. In Chupa [22], many optimizations are needed to maintain view consistency. It first generates two normal maps for a clothed human’s front and back sides, then optimizes the 3D generated mesh by the dual normal maps. There are also some human template-based works  [27, 50, 18, 48, 3, 22] that are highly dependent on human template mesh, such as SMPL [29] and SMPL-X [37], which adversely affect the diversity of generated models especially in local details. For example, it is flawed in modeling loose clothes like dresses.
@@
Based on the development of 3D human representation, numerous methods have emerged for native 3D generation. Some of these approaches [1, 34, 9, 44] utilize tri-planes [4] which is a NeRF-based representation. It is hard to generate precise geometry, which is limited by computational complexity. Apart from that, the implicit functions  [35, 36] are highly favored in several methods [46, 3, 19, 18, 24]. Alternatively, some methods use 3D data for training, gDNA [6] use implicit multi-subject forward skinning which enables learning from 3D scans of human bodies. Among all the works mentioned above, the GAN [13, 50, 19, 1, 34, 9, 20, 4, 35, 46] and Diffusion models [17, 44, 3, 22, 24] are the most popular.
Using the diffusion model for small object generation[7, 26, 8, 5] is more straightforward to combine with text control. However, the text-guided 3D human generation is more difficult due to the high-dimensional cube and lack of text-3D data pairs. Some attempts were made to tackle these issues. AvatarCLIP [18] initialize a bare mesh shape with SMPL. It continuously optimizes the mesh to match the input text by calculating the CLIP score between the text token and the mesh’s rendering image. CLIP-actor [48] does the same thing, except it designs a recommendation module to get the initial shape/motion. In this way, they reduced the need for text-3D data pairs with the help of the large text-image pre-trained model [40]. Other methods [22, 3, 44] impose textual controls on the diffusion model for generated shapes that match the input description. In previous approaches [8, 26, 7] to generating small objects, the resolution of the generated geometry is limited for human generation, such as $64^{3}$ or $128^{3}$. In addition, some methods, such as differentiable rendering-based methods, are time-consuming to infer.
@@ 3 Method
Figure 2: Overview of our framework. Joint2Human takes Gaussian noise and some conditional information as input. We first encode FOF into a latent space, where a diffusion model is trained to enable high-resolution human generation. Furthermore, to enable flexible conditional generation, we adopt condition encoders along with classifier-free guidance to enable conditional generation. Our conditional control generation strategy can support switching between different modalities. Then, we pass the results generated by the first denoising process through the high-frequency enhancer $\mathcal{D}_{h}$ and multi-view recarving strategy for fine-grained local detail generation. Additionally, the recaiving strategy is shown in the bottom subfigure.
Figure 3: The data processing flow for compact spherical embedding of 3D joints.
Our method aims to generate diverse 3D-clothed humans with global structure and local details. First, we propose a direct conditional 3D generative method with 2D diffusion models as illustrated in Fig. 2. Specifically, we utilize an image-aligned 3D representation FOF to produce 3D shapes directly. We operate the condition and generation process in the latent space (Sec. 3.1). Then, we propose the compact spherical embedding of 3D joints (Sec. 3.2) for precise pose control and diverse generation. Furthermore, a high-frequency enhancer (Sec. 3.3) and a multi-view recarving strategy (Sec. 3.4) are proposed to improve the geometry details further. We will detail all the stages of our method in the following subsections.
@@ 3.1 Latent Diffusion for Fourier Occupancy Fields
To model the distribution of Fourier Occupancy Fields, following the latent diffusion model [41], we utilize an auto-encoder to compress the high-dimensional raw data space into a lower-dimensional latent space, which encodes human shapes into a normal distribution. In detail, we adopted a VAE [23], which contains the encoder $\mathcal{E}$ and decoder $\mathcal{D}$. Given a FOF feature $x\in\,\,\mathbb{R}^{512\times 512\times 32}$, the encoder $\mathcal{E}$ encodes FOF into latent vectors $z=\mathcal{E}\left(x\right)$, where $z\in\mathbb{R}^{128\times 128\times 8}$. The decoder $\mathcal{D}$ decodes the latent vectors back to the original FOF space $\tilde{x}=\mathcal{D}\left(\mathcal{E}\left(x\right)\right)$, where $\tilde{x}\in\mathbb{R}^{512\times 512\times 32}$. We pretrained the auto-encoder with the reconstruction loss and KL-regularization loss like [41]. During training, the training data is passed through the encoder $\mathcal{E}$ in the vae module to obtain its feature in the latent space. We adopt a U-Net-like structure as our denoising model $\epsilon_{\theta}$. We follow the classical and efficient loss function proposed by Ho [17] and train the diffusion model based on the $T$ steps of noise-adding and denoising processes in the latent space. Thus, we can obtain a vector $\tilde{z}_{0}$ in the latent space, which will be sent to the decoder $\mathcal{D}$ to get FOF. Then, we can convert the FOF to an occupancy field and get an initial 3D human mesh from it with the Marching Cubes [30] algorithm.
@@ 3.2 Condition-guided Generation Mechanisms
Compact Spherical Embedding of 3D Joints. Different from previous works [22, 9, 50, 48, 18, 3] adopt the human parametric model [29, 37] for pose guidance, we find that such a strategy with FOF can lead to overfitting. Conditioning on the SMPL makes it harder for networks to learn the true data distribution, which is also mentioned in ECON [47]. To avoid the redundant information and the misleading geometric prior from SMPL, we use the human body joints as pose guidance and design a compact spherical embedding of 3D joints. As shown in Fig. 3, we first estimate the $K$ 3D joints location $J=\left\{p_{i}\in\mathbb{R}^{3}\right\}_{i=1}^{K}$ of the given human. For each joint position $p_{i}$, we extend a sphere with $p_{i}$ as the center and $r$ as the radius in 3D space. After that, according to the Sec. 3.1, we compute the FOF $c_{i}\in\mathbb{R}^{512\times 512\times 8}$ for each sphere. Therefore, the depth information is stored in the phase of trigonometric functions. We concatenate $\left\{c_{i}\right\}_{i=1}^{K}$ channel-wise with a fixed order to form the compact spherical embedding of 3D joints $con_{joint}\in\mathbb{R}^{512\times 512\times 8K}$. In this way, our approach processes joints in different channels and concatenates them to form the embedding of 3D joints, integrating full semantic and depth-wise information for precise pose control. In our setting, we use $r=10cm$ and follow the joints $K=24$ in SMPL.
@@
Pose-guided Geometry Generation. We aim to generate a human shape that fits the given pose. With the compact spherical embedding of the 3D joints, we can efficiently perform pose control in human generation. To further improve the stability of the generative model, we incorporate the 2D prior by utilizing the IUV map $con_{iuv}$ defined in DensePose [14]. During training, we first get the $con_{pose}$ by concatenating the $con_{joint}$ and $con_{iuv}$ in the channel dimension. Then, we learn a pose encoder $\varepsilon_{pose}$ to get the human pose embedding $\tau_{pose}$ and apply the pose condition $con_{pose}$ on the latent code $z$ by concatenation after passing through the conditional encoder :
$\tau_{pose}=\varepsilon_{pose}\left(con_{pose}\right),$ (1)
where the $\varepsilon_{pose}$ consists of two convolutional layers for the alignment of the feature. In the inference stage, we first get the human pose embedding $\tau_{pose}$; then we perform the denoising process of the $\epsilon_{\theta}$ to get the initial human shape.
@@
Text-guided Geometry Generation. To generate the 3D human with the guidance of text description, we proposed the image prompt strategy for our model based on a text-guided 2D human image generation model Text2Human [21]. In the training stage, we learn an additional image encoder $\varepsilon_{i}$ to get the human image embedding $\tau_{img}$, then apply it to the training of the diffusion model by concatenation. In the inference stage, with the input text, we first infer the text-to-image model $\mathcal{T}$ and image encoder $\varepsilon_{i}$ to get the embedding $\tau_{img}$, then we sample the 3D human shape from $\epsilon_{\theta}$ with all these conditions. We provide additional text-guided generation results in the supplementary materials.
3.3 High-frequency Enhancer
For efficient human representation and modeling of data distribution, we only save low-frequency terms while taking the FOF as a 3D representation in  Sec. 3.1. However, some abandoned high-frequency information is essential for human shape details. Therefore, based on the generation of the low-frequency FOF feature, we propose the High-Frequency enhancer to recover these missing details $C_{h}$, as shown in Fig. 2. In detail, we learn a reference-based decoding network $\mathcal{D}_{h}$ based on the latent space and decoder from the auto-decoder claimed in Sec. 3.1 to predict high-frequency terms $\hat{C}_{h}$:
$\hat{C}_{h}=\mathcal{D}_{h}\left(\mathcal{D}_{l}\left(z_{0}\right),z_{0}\right),$ (2)
To enhance the perception of latent-space contextual information, we introduce additional skip connections for each layer in $\mathcal{D}_{h}$, which is used to fuse the feature map $\mathcal{D}_{l}\left(z_{0}\right)$ from the relevant layers in the decoder $\mathcal{D}_{l}$. When training, we calculate the MSE loss between the predicted terms ${C}_{h}$ and high-frequency ground truth ${C}_{h}$ as the supervision to optimize $\mathcal{D}_{h}$,the loss function $\mathcal{L}_{h}$ is formulated as:
$\mathcal{L}_{h}=\frac{1}{n}\sum_{i=1}^{n}{\left(\hat{C}_{h}\left(x_{i},y_{i}\right)-C_{h}\left(x_{i},y_{i}\right)\right)^{2}},$ (3)
where $n$ is the number of all pixels in ${C}_{h}$. So, in the inference stage, we can estimate the high-frequency feature based on the known low-frequency feature to enhance the geometric details.